A remotely located air sampling station can be powered by solar cells or by running an electric line to the site and using conventional power. Solar cells will cost $8,000 to install and will have a useful life of 4 years with no salvage value. Annual costs for inspection, cleaning, etc. are expected to be $1,700. A new power line will cost $15,000 to install, with power costs expected to be $1,400 per year. Since the air sampling project will end in 4 years, the salvage value of the line is considered to be zero. At an interest rate of 10% per year, which alternative should be selected on the basis of a future worth analysis?

The future worth of solar cells is $______ and that of electric line is $________ .
Answer #1

Solar cells

Power line

Initial Cost = 8,000

Life = 4 years

Salvage Value = 0

Annual cost = 1,700

Interest rate = 10%

Future Worth = 8,000 (F/P, 10%, 4) + 1,700 (F/A, 10%, 4)

Future Worth = 8,000 (1.4641) + 1,700 (4.6410)

Future Worth = 19,602.5

Initial Cost = 15,000

Life = 4 years

Salvage Value = 0

Annual power cost = 1,400

Interest rate = 10%

Future Worth = 15,000 (F/P, 10%, 4) + 1,400 (F/A, 10%, 4)

Future Worth = 15,000 (1.4641) + 1,400 (4.6410)

Future Worth = 28,458.9

Which alternative should be selected on the basis of a future worth analysis?

Alternative SOLAR CELLS should be selected as it has less cost than the power line.

Select – SOLAR CELLS